# Stable-Hair: Real-World Hair Transfer via Diffusion Model

### v1

@@ Introduction
As illustrated in the figure 6, we conducted qualitative comparison experiments on a variety of hairstyles. Overall, our method significantly outperforms other approaches in terms of the refinement and completeness of hairstyle transfer, while also maintaining the structural and identity consistency of the source image to a great extent. Among the methods we compared, Barbershop frequently exhibits simple hair pasting issues and often result in unnatural integration of hair and face (as seen in the third and fourth rows). Style Your Hair and Hair Fast GAN performs rather coarsely in hairstyle transfer, often neglecting hair details and showing color inconsistencies, especially with colored hair. Hairclip and Hairclip v2 demonstrate the weakest ability in hairstyle transfer, struggling to accurately transfer the reference hairstyle. In contrast, our method consistently exhibits robust and stable transfer capabilities across different hairstyle styles and colors.
Quantitative Comparison
[…]
Barbershop 46.178 29.324 0.579 0.760
HairFastGAN 36.205 29.348 0.620 0.762
SYH 41.543 29.517 0.621 0.712
Hairclip 44.359 28.056 0.620 0.697
HairclipV2 37.456 30.619 0.637 0.769
Ours 33.653 29.555 0.640 0.771
Figure 7: Ablation study results of different training and inference settings.
The experiment uses the CelebA-HQ datasets as experimental data, 2500 face images are randomly selected as input from the CelebA-HQ dataset with an equal number of reference images from the remaining Celeb-HQ dataset. Table 1 shows our quantitative evaluation across different methods. Overall, our method surpasses previous approaches on most metrics. Specifically, the higher SSIM and IDS scores indicate that Stable-Hair effectively preserves the content and identity of the source image while transferring the hairstyle, demonstrating our method’s superiority in maintaining structural and content consistency. Additionally, the superior FID score suggests that our method produces high-fidelity and highly realistic results after the transfer. Our method ranks second in terms of the PSNR metric. Although it is slightly inferior to Hairclip v2 in this specific measure, our method still demonstrates high image fidelity and quality when considering other comprehensive metrics.
@@
Table 2: User study on hair transfer. Accuracy denotes the accuracy for hair transfer, Preservation indicates the ability to preserve irrelevant regions and Naturalness denotes the visual realism of the generated image.
@@
Figure 8: The visual results of cross-domain transfer demonstrate the robustness of our method.
Ablation study.
To thoroughly investigate the role of each module in our method, we conducted systematic ablation studies by sequentially removing individual components. As illustrated in figure 7, it is evident that the models trained with pixel-conditioned input using ControlNet often exhibit color discrepancies (as seen in the first and third columns of the results), leading to inconsistencies between the source and target images. By employing our proposed Latent ControlNet, which maps the hair removal and transfer process from pixel space to latent space, we effectively eliminate these color inconsistencies, significantly enhancing content preservation. Additionally, the bald images generated in our Stage 1 output appear very natural, allowing for a seamless transition into Stage 2. This results in highly satisfactory transfer outcomes.
Cross-domain transfer results.
As illustrated in figure 8, the robustness of our method enables the transfer of hairstyles across diverse domains, a capability that previous approaches were unable to achieve. This demonstrates the significant advancements our method offers in the realm of hairstyle transfer.
@@ Conclusions
In this paper, we introduce Stable-Hair, the first framework to tackle hairstyle transfer using diffusion techniques. This method marks a significant advancement, achieving stable and fine-grained real-world hairstyle transfers previously unattainable. Stable-Hair features a two-stage pipeline. The first stage uses a Bald Converter to transform the source image into a bald proxy image. The second stage comprises a Hair Extractor, a Latent IdentityNet, and Hair Cross-Attention Layers to accurately transfer the target hairstyle to the bald image. The Hair Extractor encodes reference images with the desired hairstyles, while the Latent IdentityNet preserves identity content and background consistency. The Hair Cross-Attention Layers in the U-Net ensure precise and high-fidelity hairstyle transfer. Extensive experiments demonstrate that Stable-Hair achieves commercial-grade hairstyle transfer capabilities, setting a new standard in the field.
@@
Diffusion Model (DM) (Ho, Jain, and Abbeel 2020) belongs to the category of generative models that denoise from a Gaussian prior $\mathbf{x_{T}}$ to target data distribution $\mathbf{x_{0}}$ by means of an iterative denoising procedure. Latent Diffusion Model (LDM) (Rombach et al. 2022) is proposed to model image representations in the autoencoder’s latent space. LDM significantly speeds up the sampling process and facilitates text-to-image generation by incorporating additional text conditions. The LDM loss is:
@@
Figure 9: Limitation of our method.
@@
More Visual Results
@@
Broader Impact
Limitations
Due to the limitation of the training data, our method may inadvertently transfer certain hair accessories along with the makeup to the source image (As shown in Fig 9), which may not be desirable in some scenarios. We plan to address this limitation in future work to achieve more controlled and precise hair transfer.

### v2

@@ Introduction
As illustrated in Fig. 4, we conducted qualitative comparison experiments on a variety of hairstyles. Overall, our method significantly outperforms other approaches in terms of the refinement and completeness of hairstyle transfer, while also maintaining the background and identity consistency of the source image to a great extent. Among the methods we compared, Barbershop tends to produce chaotic results for complex hairstyle transfers, as observed in the second and fifth rows. Style Your Hair and HairFastGAN perform relatively coarsely in hairstyle transfer, often neglecting the fine details of hair texture and color. Hairclip and Hairclip v2 demonstrate the weakest ability in hairstyle transfer, struggling to accurately transfer the reference hairstyle. In contrast, our method consistently exhibits robust and stable transfer capabilities across different hairstyle styles and colors. Furthermore, We also compared the hair removal capabilities of our Bald Converter with HairMapper in Fig. 5. Our Bald Converter demonstrates robust and reliable performance, effectively transforming source images across various challenging scenarios. In contrast, HairMapper frequently struggles in these contexts, leading to inconsistencies in both identity and background.
Quantitative Comparison.
Figure 5: Visual comparison of hair removal using HairMapper. Our bald converter demonstrates robust performance, effectively converting source images across a wide range of poses, half-body shots, and even animated characters. In contrast, HairMapper struggles with these diverse scenarios, failing to maintain identity and background consistency.
The experiment uses the CelebA-HQ dataset (Karras et al. 2017) as experimental data, 2500 face images are randomly selected as input from the CelebA-HQ dataset with an equal number of reference images from the remaining Celeb-HQ dataset. Table 1 shows our quantitative evaluation across different methods. Overall, our method outperforms previous approaches on the majority of metrics, attaining either the top or second position across all evaluated criteria. Specifically, the highest SSIM and IDS scores, along with a notable PSNR ranking, demonstrate that Stable-Hair effectively preserves both the background and identity of the source image during hairstyle transfer. Additionally, the lowest FID score indicates that our method achieves superior visual fidelity and realism in the transferred hairstyles. Achieving the highest CLIP-I score indicates that our method excels at aligning the transferred hairstyle with the reference image, showing strong performance in maintaining the desired style.
@@
Table 2: Quantitative ablation of different design options. (‘pixel’ refers to the use of the original ControlNet, while ’latent’ denotes the use of the Latent ControlNet.)
[…]
Ours (latent) 0.434 35.128 30.980 0.680 0.779
Ours (pixel) 0.423 39.236 29.317 0.668 0.770
Figure 6: Visual ablation study results comparing different design options. The Latent ControlNet clearly maintains color consistency, whereas the original ControlNet introduces color discrepancies at each stage.
Ablation Study.
To thoroughly investigate the role of each module in our method, we conducted systematic ablation. As illustrated in Fig. 6, it is evident that models trained with pixel-conditioned input using ControlNet often exhibit color discrepancies (as seen in the first and third columns of the results), leading to inconsistencies between the source and target images. Using our proposed Latent ControlNet, which maps the hair removal and transfer process from pixel space to latent space, we effectively eliminate these color inconsistencies, significantly enhancing content preservation. The results of the quantitative ablation are shown in Table 4. Our method consistently outperforms across all metrics, highlighting the superior transfer capabilities and the enhanced preservation of identity and background achieved by the Latent ControlNet.
To further investigate the impact of video data on our method, we conducted training using only image data while keeping the same parameter settings. The results of the visual ablation study, as shown in Fig. 7, clearly indicate that when trained solely with image data, our method struggles to handle significant variations in facial poses from the source images, leading to a collage-like effect with less natural transitions. In contrast, when trained with a mix of video data, our method demonstrates robustness to changes in facial pose. See the supplements for additional ablation studies.
Table 3: User study on hair transfer. Accuracy denotes the accuracy for hair transfer, Preservation indicates the ability to preserve irrelevant regions and Naturalness denotes the visual realism of the generated image. Our method achieves the best results in all three categories.
@@
Figure 7: Visual ablation results of different training data. When trained exclusively on image data, the model often resorts to a simplistic copy-paste of the hairstyle. However, by integrating video data into the training process, our method gains pose robustness, enabling adaptive and natural hair transfer based on facial orientation.
@@ Conclusions
In this paper, we introduce Stable-Hair, the first framework to tackle hairstyle transfer using diffusion techniques. This method marks a significant advancement, achieving stable and fine-grained real-world hairstyle transfers previously unattainable. Stable-Hair features a two-stage pipeline. The first stage uses a Bald Converter to transform the source image into a bald proxy image. Following this step, we designed an automated data collection pipeline that leverages the Bald Converter, an inpainting model, and ChatGPT to gather a diverse and large-scale dataset of triplets for training the second stage. In the second stage, a Hair Extractor and a Latent IdentityNet are utilized to accurately and robustly transfer the target hairstyle onto the bald image. Extensive experiments demonstrate that Stable-Hair achieves commercial-grade hairstyle transfer capabilities, setting a new standard in the field.
@@
Diffusion Model (DM) (Ho, Jain, and Abbeel 2020) belongs to the category of generative models that denoise from a Gaussian prior $\mathbf{x_{T}}$ to the target data distribution $\mathbf{x_{0}}$ using an iterative denoising procedure. Latent Diffusion Model (LDM) (Rombach et al. 2022) is proposed to model image representations in the autoencoder’s latent space. LDM significantly speeds up the sampling process and facilitates text-to-image generation by incorporating additional text conditions. The LDM loss is as follows:
@@
Source Data
We utilize these three datasets to collect our training data.
  • FFHQ dataset subset (Karras, Laine, and Aila 2019): We randomly selected 20,000 images from the FFHQ dataset as the original image data for our training.
  • CelebV-HQ (Zhu et al. 2022a): Large-scale, high-quality video data containing 35,666 clips involving 15,653 identities and 83 manually labeled facial attributes.
Table 4: Quantitative ablation of different training data.
[…]
Ours (image+video) 0.434 35.128 30.980 0.680 0.779
Ours (only image) 0.431 33.653 30.541 0.676 0.771
Quantitative Ablation of Training Data
In this section, we further examine the quantitative impact of incorporating video data into the training set on the performance of hairstyle transfer. As shown in Table 4, when the training data consists solely of images, the FID metric shows improvement, while other metrics exhibit slight declines compared to the case where the full dataset is used. This indicates that while there is a slight decline in fidelity and overall image generation quality when video data is incorporated into the training, the model’s ability to transfer hairstyles and maintain identity and background consistency is enhanced
Why Use Bald Converter?
First, removing the original hairstyle eliminates the need for the model to blend the new hairstyle with any remnants of the old one, reducing visual artifacts and resulting in a more natural and visually appealing transfer. Secondly, by standardizing the source images to a bald state, the model can better focus on facial features and structure, enhancing the accuracy and realism of the hairstyle transfer. Additionally, this preprocessing step helps to reduce variability in the training data, improving the model’s generalization capability across different hairstyles. Lastly, converting to a bald state simplifies the alignment process, ensuring that the new hairstyle seamlessly integrates with the subject’s face, even in cases of diverse facial poses.
To further demonstrate the importance of standardizing the source image’s hairstyle to a bald state in Stage 1, we conducted a set of comparative experiments. In these experiments, rather than standardizing to a bald state as in Stage 1, we randomly edited the source image’s hairstyle using an inpainting model and then trained a hair transfer model based on these edited images. The training framework and data references are illustrated in the figure 9, with all training settings consistent with those described in our experiments section. The visual comparison results are also shown in the figure 10. As evident from the results, our model, which standardizes the hairstyle to a bald state, achieves high-fidelity and precise hair transfer, whereas the model trained on randomly edited hairstyles fails to exhibit any effective hairstyle transfer capability.
Additionally, our method achieves hair removal without requiring image alignment or cropping, ensuring that the entire hairstyle is accurately removed and no residual hair remains. This guarantees a cleaner, more consistent starting point for subsequent hairstyle transfer, leading to improved visual fidelity and more realistic results. As shown in Fig. 12, our method is capable of removing hair from images with various poses, half-body shots, and even cartoon or animated characters.
@@
More Transfer Results
Figure 8: Limitation of our method.
Limitations
Due to the limitation of the training data, our method may inadvertently transfer certain hair accessories to the source image (as shown in Fig. 8), which may not be desirable in some scenarios. We plan to address this limitation in future work to achieve more controlled and precise hair transfer.
Broader Impact
Figure 9: The training framework without the bald converter. (i.e., the source data is not standardized to the bald state)
Figure 10: Visual comparison of the results under different data strategies reveals significant differences. When the bald converter is not applied (i.e., the source data is not standardized to the bald state), the model fails to learn how to transfer hairstyles during training and instead generates random hairstyles. However, when the data is standardized to the bald state, the model incrementally learns to transfer high-fidelity and high-precision hairstyles.
@@
Figure 12: More visual results of hair removal.